Grade 11Mechanics

a projectile has the same range rwhen the maximum ht attained by it is either div or h2.find the relation between div,h2 and r

maazhussain
9 Years agoGrade 11

1 Answer

Samyak
9 Years ago
When range is same then angle of projection is complementary.H1= u^2sin^2x/2gH2=u^2cos^2x/2gH1^2+H2^2 is u^2/2gTherefore h1^2+h2^2=rsin2x/2x is angle of projection.Please approve my answer.